find 3 rational numbers lying between 0 and 0.1. find 20 rational numbers between 0and0.1.give a method determine any number of rational numbers between 0 and 0.1
 

Grade:10

2 Answers

Arun
6 years ago
For rational number between 0 and 0.1
3 rational number = 0.05, 0.051, 0.052
10 rational numbers = 0.05, 0.051, 0.052,0.053,.....so on
Hencce you can write any number between two rational number
Subhash Kumar Mehta
6 years ago
we have a=0,b=0.1,and n=3let no. be a+d,a+2d,a+3dwhere d=(b-a)/(n+1)so d=(0.1-0)/(3+1)=1/40so nos. are 0+(1/40),0+2*(1/40),0+3*(1/40)=1/40,1/20,3/40so we can find n number between a and b as above and also 20 numbers by this method